OpenStudy (anonymous):

Can someone please help me with this problem. I don't get it. O.o Which of the following points satisfies the inequality 2x - 3y < 1? A. (-2, 1) B. (1/2, 0) C. (2, -1)

OpenStudy (anonymous):

Just plug each point in and see if the end result is true. A. (-2,1) 2x - 3y < 1 = 2(-2) - 3(1) < 1 = -4 - 3 < 1 = -7 < 1 Is this true? Yes, it is, because -7 is less than one. B. (1/2, 0) 2x - 3y < 1 = 2(1/2) - 3(0) < 1 = 1 - 0 < 1 = 1 < 1 Is this true? No, because 1 is not less than 1. C. (2, -1) 2x - 3y < 1 = 2(2) - 3(-1) < 1 = 4 - (-3) < 1 = 4 + 3 < 1 = 7 < 1 Is this true? No, because 7 is not less than 1. So your answer? A, because it is the only one that satisfies the inequality. (: Does this clarify things?

OpenStudy (anonymous):

ohhh okay i see thank you!!!